OpenStudy (anonymous):

In 1987, the Australian humpback whale population was 350 and has increased at a rate of about 14% each year since then. Write a function to model population growth. Use a graph to predict when the population will reach 20, 000.

OpenStudy (anonymous):

Option A - 30.9. :P :D

OpenStudy (anonymous):

A.P(t) = 15,500(0.97)t; the population will drop below 8,000 in 21.7 yr. B.P(t) = 15,500(0.97)t; the population will drop below 8,000 in 27 yr. C.P(t) = 15,500(1.03)t; the population will drop below 8,000 in 21.7 yr. D.P(t) = 15,500(1.03)t; the population will drop below 8,000 in 27 yr.

OpenStudy (anonymous):

@bajji

OpenStudy (anonymous):

a)P(t) = 350(1.14)t; the population will reach 20,000 in about 30.9 yr. b)P(t) = 350(0.86)t; the population will reach 20,000 in about 19.5 yr. c)P(t) = 350(0.86)t; the population will reach 20,000 in about 39.1 yr.wwwwww d)P(t) = 350(1.14)t; the population will reach 20,000 in about 25.6 yr. Not these ones?? O_o

OpenStudy (anonymous):

Check properly ! The options provided by you doesn't match your question!!